# Trained Random Forests Completely Reveal your Dataset

### v1

@@ 7.1 Experimental Setup
We rely on three popular datasets for binary classification in our experiments. We discretize each dataset’s numerical attributes, and one-hot encode the categorical ones. To keep a reasonably small number of features, we remove some attributes with the smallest support [^2] Our binarized versions of these datasets are available in the supplementary material and will be available on our online repository upon publication.. First, the COMPAS dataset (analyzed by Angwin et al. (2016)) gathers records about criminal offenders in the Broward County of Florida collected from 2013 and 2014, with the task being recidivism prediction. Our preprocessed version includes $7,206$ examples described by $15$ binary attributes. Second, the UCI Adult Income dataset (Dua & Graff, 2017) contains data regarding the 1994 U.S. census to predict whether a person earns more than \$50K/year. After preprocessing, our dataset includes $48,842$ examples and $20$ binary features. Finally, we use the Default of Credit Card Client dataset (Yeh & hui Lien, 2009), to predict whether a person will default in payment (the next time they use their credit card). Our preprocessed version includes $29,986$ examples and $22$ binary attributes.
@@ 7.2 Results
The results of our experiments are reported in Figure 2 for all three datasets, with or without the use of bagging to train the target RFs. More precisely, we plot the average reconstruction error as a function of the number of trees $\lvert\mathcal{T}\rvert$, for several values of the trees’ maximum depth $d_{max}$. We observe several trends that are consistent across all three datasets. In all cases, as expected, increasing the trees’ depth or the number of trees in the forest decreases the reconstruction error as it provides more information regarding the training data. When bagging is not used to train the RFs, the reconstruction error reaches 00 in all cases for the deepest forests (recall that the default parameters of the scikit-learn library is no maximum depth constraint). This is not the case when using bagging. In such cases, the reconstruction error reaches a threshold and stops improving even for larger forests. In Appendix B, we further investigate the effect of bagging on protecting the training data against reconstruction attacks and conduct additional experiments. Our main finding is that this performance drop precisely comes from the difficulty of guessing how many times each example went through each tree: bagging intrinsically provides a form of protection regarding the training data. This is consistent with theoretical results stating that bagging provides (weak) differential privacy guarantees (Liu et al., 2021b).
We report in Tables 2 and 3 (respectively for a fixed number of trees $\lvert\mathcal{T}\rvert=100$ and no fixed maximum depth, both corresponding to scikit-learn’s default values) the average run times for the reconstruction model without bagging, along with the number of times (#Runs) the solver was unable to find a feasible solution before timeout. When bagging is used, most runs attain the time limit and return a solution but cannot prove optimality. Run times in this context are not informative, so we only report the number of times the solver did not find any feasible solution before the time limit. Note that the few runs that did not produce a feasible solution are excluded from Figure 2. We observe from Table 2 that the formulation without bagging efficiently handles the problems that are under-constrained (shallow trees) or over-constrained (deep trees). Intermediate cases seem to require more computational effort, and in a few cases, the solver did not find a feasible solution. When using bagging, the size of the models seems to matter the most, as the solver only failed to find feasible solutions with the deepest forests. The same observation holds from Table 3, as the only runs for which the solver did not find a feasible solution are those with the largest numbers of trees. When not using bagging, the solution times scale approximately linearly with the number of trees.
As discussed in Section 3, most works in the reconstruction attacks literature only target reconstructing part of the dataset attributes (generally, a single one), assuming the others are publicly known. In the Appendix D, we perform complementary experiments on such partial reconstruction. The results show that our approach successfully leverages knowledge of part of the dataset attributes, which results in lower error rates for the other ones.
@@ 8 Discussion and Conclusions
This study has shown that the structure of a trained RF can be exploited to reconstruct most (if not all) of its training data. It introduced an attack leveraging mathematical programming tools to encode the structure of an RF and rely on a general-purpose CP solver to find the most likely reconstructions of the training data. Due to the high redundancy of RFs built using off-the-shelf ML libraries along with their default parameters, the resulting problem is often strongly constrained, resulting in a high reconstruction rate. While theoretical $\mathcal{NP}$-completeness theorems indicate that such an attack may not be computationally tractable at scale, the tremendous progress in CP/MILP solvers has made it practical to solve larger and larger problems over time. Therefore, it may just be a question of time until data breaches happen for large datasets.
The fact that the proposed framework is based on mathematical programming techniques opens the door to many promising research perspectives. For instance, the approach could be extended to handle other types of attributes (numerical, categorical) without the need for feature binarization. Performance improvements could also be achieved through different problem reformulations or additional valid inequalities. Notably, one could leverage the information gain criterion used to select the splits while building the decision tree to eliminate combinations of attributes’ values leading to different splits.
Our framework can also be used as a building block for other types of inference attacks, such as membership inference or property inference. Furthermore, we considered canonical RFs trained without privacy-preserving techniques, representing most of what popular libraries do by default. Investigating the effectiveness of common privacy-preserving mechanisms, such as the widely used differential privacy (Dwork et al., 2014), would bring additional insights. Though this may lead to difficult models, the proposed CP (or MILP) formulations could be extended to infer the noise added by the protection mechanisms on the released per-node counts (Fletcher & Islam, 2019; Dinur & Nissim, 2003). Finally, another interesting direction is to apply the proposed methodology to other types of ensembles and ML models.
@@
Appendix A Mixed-Integer Linear Programming Formulation
@@
A.1 Model Formulation (Without Bagging)
@@
A.2 Empirical Evaluation
@@
Figure 3: Average reconstruction error as a function of the number of trees $\lvert\mathcal{T}\rvert$ within the attacked forest $\mathcal{T}$, for different maximum depth values $d_{max}$ and for the random baseline. For the experiments on the COMPAS dataset, not using bagging, we report the results obtained using either the CP model (Section 6) or the MILP one (Section A.1).
@@
The results are reported in Figure 3, and their run times are compared in Table 4. Note that the results for the CP model are those presented in Figure 1(a), repeated here to ease comparison. Comparing the different curves (which correspond to different maximum depth constraints) between Figures 2(a) and 2(b), we see that both approaches successfully solve the dataset reconstruction problem on COMPAS without the use of bagging to train the target random forests. Intuitively, the two feasibility models encode the same information, and define the same set of feasible reconstructions. Because they use different techniques to represent and explore it, they may end up with different reconstructions, but there is no a priori reason for one to outperform the other systematically, and as observed in our experiments, their reconstruction performances are generally similar.
Appendix B The Impact of Bagging on Data Protection
  • $\forall t\in\mathcal{T},\forall k\in\{1..N\},\forall v\in\mathcal{V}_{t}^{L}:$ if $\textbf{if }\sum\limits_{c\in\mathcal{C}}y_{tvkc}\geq 1\textbf{ then}\quad\left(\bigwedge\limits_{i\in\Phi_{v}}{x_{ki}=1}\right)\land\left(\bigwedge\limits_{i\in\overline{\Phi_{v}}}{x_{ki}=0}\right)$
Let $\phi^{y}_{tk}:=\left\{v\in\mathcal{V}_{t}^{L}:\sum\limits_{c\in\mathcal{C}}y_{tvkc}\geq 1\right\}$ be the (possibly empty) set of leaves of tree $t$ for which example $k$ has been used. While the leaf-consistency constraints fix all attributes $i$ in $\Phi_{v}\cup\overline{\Phi_{v}}$ for $v\in\phi^{y}_{tk}$, any feature that does not appear in any such sets (call them free attributes) can be arbitrarily set without changing the likelihood of the solution. And so, the fact that a free attribute is guessed correctly can be attributed to luck and should not be seen as a positive aspect of the CP model.
Formally, the fixed attributes for example $k\in\{1..N\}$ are
$\mathcal{F}_{k}:=\bigcup\limits_{t\in\mathcal{T}}\bigcup\limits_{v\in\phi^{y}_{tk}}\left(\Phi_{v}\cup\overline{\Phi_{v}}\right)$
$x_{ki}^{y}:=\begin{cases}1,&\mbox{if }i\in\mathcal{F}_{k}\cap\Phi_{v}\mbox{ for some }t\in\mathcal{T},v\in\phi^{y}_{tk}\\ 0,&\mbox{if }i\in\mathcal{F}_{k}\cap\overline{\Phi_{v}}\mbox{ for some }t\in\mathcal{T},v\in\phi^{y}_{tk}\\ 1-\mathbf{x}_{ki},&\mbox{ otherwise}\end{cases}$
Our benchmark experiment can now be described as follows:
  • Output the set of solutions $\{x_{k}^{y}\}_{k=1}^{N}$.
The results of the benchmark experiments for the three considered datasets are shown in Figures 3(b), 3(d) and 3(f). The results without bagging are also repeated in Figures 3(a), 3(c) and 3(e) (from Figures 1(a), 1(c) and 1(e)) for reference and easy comparison.
(a) COMPAS dataset, bagging not used
(b) COMPAS dataset, benchmark runs
(c) UCI Adult Income dataset, bagging not used
(d) UCI Adult Income dataset, benchmark runs
(e) Default of Credit Card Client dataset, bagging not used
(f) Default of Credit Card Client dataset, benchmark runs
@@ Appendix C Implementation details for the CP model
if $\textbf{if }\sum\limits_{c\in\mathcal{C}}y_{tvkc}\geq 1\textbf{ then}\quad\left(\bigwedge\limits_{i\in\Phi_{v}}{x_{ki}=1}\right)\land\left(\bigwedge\limits_{i\in\overline{\Phi_{v}}}{x_{ki}=0}\right)$
@@
  • $\forall t\in\mathcal{T},\forall k\in\{1..N\},\forall v\in\mathcal{V}_{t}^{L}:$ if $\textbf{if }w_{tvk}=1\textbf{ then}\quad\left(\bigwedge\limits_{i\in\Phi_{v}}{x_{ki}=1}\right)\land\left(\bigwedge\limits_{i\in\overline{\Phi_{v}}}{x_{ki}=0}\right)$

### v2

@@ 7.1 Experimental Setup
We rely on three popular datasets for binary classification in our experiments. We discretize each dataset’s numerical attributes and one-hot encode the categorical ones. To keep a reasonably small number of features, we remove some attributes with the smallest support [^1] Our binarized versions of these datasets are available in the supplementary material and will be available on our online repository upon publication.. First, the COMPAS dataset (analyzed by Angwin et al. 2016) gathers records about criminal offenders in the Broward County of Florida collected from 2013 and 2014, with the task being recidivism prediction. Our preprocessed version includes $7{,}206$ examples described by $15$ binary attributes. Second, the UCI Adult Income dataset (Dua & Graff, 2017) contains data regarding the 1994 U.S. census to predict whether a person earns more than \$50K/year. After preprocessing, our dataset includes $48{,}842$ examples and $20$ binary features. Finally, we use the Default of Credit Card Client dataset (Yeh & hui Lien, 2009), to predict whether a person will default in payment (the next time they use their credit card). Our preprocessed version includes $29{,}986$ examples and $22$ binary attributes.
@@ 7.2 Results
The results of our experiments are reported in Figure 2 for all three datasets, with or without the use of bagging to train the target RFs. More precisely, we plot the average reconstruction error as a function of the number of trees $\lvert\mathcal{T}\rvert$, for several values of the trees’ maximum depth $d_{max}$. We observe several trends that are consistent across all three datasets. In all cases, as expected, increasing the trees’ depth or the number of trees in the forest decreases the reconstruction error as it provides more information regarding the training data. When bagging is not used to train the RFs, the reconstruction error reaches 00 in all cases for the deepest forests (recall that the default parameters of the scikit-learn library is no maximum depth constraint). This is not the case when using bagging. In such cases, the reconstruction error reaches a threshold and stops improving even for larger forests. In Appendix E, we further investigate the effect of bagging on protecting the training data against reconstruction attacks and conduct additional experiments. Our main finding is that this performance drop precisely comes from the difficulty of guessing how many times each example went through each tree: bagging intrinsically provides a form of protection regarding the training data. This is consistent with theoretical results stating that bagging provides (weak) differential privacy guarantees (Liu et al., 2021b).
@@
We report in Tables 2 and 3 (respectively for a fixed number of trees $\lvert\mathcal{T}\rvert=100$ and no fixed maximum depth, both corresponding to scikit-learn’s default values) the average run times for the reconstruction model without bagging, along with the number of times (#Runs) the solver was unable to find a feasible solution before timeout. When bagging is used, most runs attain the time limit and return a solution but cannot prove optimality. Run times in this context are not informative, so we only report the number of times the solver did not find any feasible solution before the time limit. Note that the few runs that did not produce a feasible solution are excluded from Figure 2. We observe from Table 2 that the formulation without bagging efficiently handles the problems that are under-constrained (shallow trees) or over-constrained (deep trees). Intermediate cases seem to require more computational effort, and in a few cases, the solver did not find a feasible solution. When using bagging, the size of the models seems to matter the most, as the solver only failed to find feasible solutions with the deepest forests. The same observation holds from Table 3, as the only runs for which the solver did not find a feasible solution are those with the largest numbers of trees. When not using bagging, the solution times scale approximately linearly with the number of trees. We report in Appendix F additional experiments regarding our method’s scalability with respect to the number of training examples $N$. The results demonstrate its ability to reconstruct considerably larger datasets, with the reconstruction error remaining very small. Furthermore, while the size of the CP model’s search space increases exponentially with the number of reconstructed training examples $N$, in practice, reconstruction time increases polynomially (approximately quadratic or sub-quadratic) with $N$.
As discussed in Section 3, most works in the reconstruction attacks literature only target reconstructing part of the dataset attributes (generally, a single one), assuming the others are publicly known. In Appendix G, we perform complementary experiments on such partial reconstruction. The results show that our approach successfully leverages knowledge of part of the dataset attributes, which results in lower error rates for the other ones.
@@ 8 Discussion and Conclusions
This study has shown that the structure of a trained RF can be exploited to reconstruct most (if not all) of its training data. It introduced a new paradigm of attack, leveraging mathematical programming tools to encode the structure of an RF and relying on a general-purpose CP solver to find the most likely reconstructions of the training data. Due to the high redundancy of RFs built using off-the-shelf ML libraries with their default parameters, the resulting problem is often strongly constrained, resulting in a high reconstruction rate. While theoretical $\mathcal{NP}$-completeness theorems indicate that such an attack may not be computationally tractable at scale, the tremendous progress in CP/MILP solvers has made it practical to solve larger and larger problems over time. Therefore, it may just be a question of time until data breaches happen for large datasets.
The fact that the proposed framework is based on mathematical programming techniques opens the door to many promising research perspectives. The approach could be tested on various types of attributes (numerical, categorical) without the need for feature binarization. Performance improvements could also be achieved through different problem reformulations or additional valid inequalities. Notably, one could leverage the information gain criterion used to select the splits while building the decision tree to eliminate combinations of attributes’ values leading to different splits.
Our framework can also be used as a building block for other types of inference attacks, such as membership inference or property inference. Furthermore, we considered canonical RFs trained without privacy-preserving techniques, representing most of what popular libraries do by default. Investigating the effectiveness of common privacy-preserving mechanisms, such as the widely used differential privacy (Dwork et al., 2014), would bring additional insights. Though this may lead to difficult models, the proposed CP (or MILP) formulations could be extended to infer the noise added by the protection mechanisms on the released per-node counts (Fletcher & Islam, 2019; Dinur & Nissim, 2003). On the same line, adapting the formulation to work without the knowledge of the per-leaf per-class counts (hence only supposing that each leaf contains at least one example from the predicted class), or considering other gray-box setups, are interesting directions. Finally, another interesting direction is to apply the proposed methodology to other types of ensembles and ML models.
@@
Appendix A Proof of NP-Hardness (Theorem 5.1)
The description of the construction of the instance of (DRP) can be found in the main text. We start by providing an example to clarify the construction.
Consider the following 3-SAT instance with $|L|=3$ clauses and $V=4$ variables:
$(u_{1}\vee\bar{u_{2}}\vee\bar{u_{3}})\wedge(u_{1}\vee u_{2}\vee u_{4})\wedge(\bar{u_{2}}\vee\bar{u_{3}}\vee\bar{u_{4}})$ (1)
Figure 3 shows the constructed (DRP) instance arising from it.
$f_{5}$$f_{1}$$f_{2}$$f_{2}$$f_{3}$$f_{3}$$f_{3}$$f_{3}$12121212111111111111001111111111111111
(a) Tree corresponding to clause 1
$f_{6}$$f_{1}$$f_{2}$$f_{2}$$f_{4}$$f_{4}$$f_{4}$$f_{4}$12121212001111111111111111111111111111
(b) Tree corresponding to clause 2
$f_{7}$$f_{2}$$f_{3}$$f_{3}$$f_{4}$$f_{4}$$f_{4}$$f_{4}$12121212111111111111111111111111111100
(c) Tree corresponding to clause 3
$f_{5}$$f_{6}$$f_{7}$121212126666001111
(d) Auxiliary tree
Figure 3: Example of (DRP) instance originating from (1). The left branches correspond to setting the feature to 0. The right ones set the feature to 1. The numbers below are the $n_{lvc}$ values.
Theorem 5.1 follows from Claims A.1 and A.2.
Claim A.1.
If the 3-SAT instance is a YES instance, then (DRP) is feasible.
Proof.
Suppose that there is an assignment of values to the $V$ variables of 3-SAT so that all clauses $L$ are satisfied.
Then for each clause $l\in L$, pick the leaf $v$ of the perfect binary subtree that corresponds to the assignment of variables in 3-SAT. We set $y_{lv(6|L|+1)}=1$ and $y_{lv^{\prime}(6|L|+1)}=0$ for all $v^{\prime}\in\mathcal{V}_{l}^{L}$, $v^{\prime}\neq v$. We set the first $V$ attributes of $x_{6|L|+1}$ to match the assignment of the $V$ variables that satisfy 3-SAT. We set the attributes $V+1,\ldots,V+|L|$ to 1.
For the remaining examples, we set the attributes $V+l$ of examples $6(l-1)+1,\ldots,6l$ to 1, and all other attributes in $V+1,\ldots,V+|L|$ to 0. This is done for all $l=1,\ldots,|L|$. In addition, we set the remaining attributes to match one of the leaves that have $n_{lvc}=1$ but don’t correspond to the assignment of variables in 3-SAT and set the corresponding $y_{lvk}$ to 1.
This can be easily checked to be feasible for (DRP).
For example, in Figure 3, if the 3-SAT assignment is $u_{1}=\textsc{True}$, $u_{2}=u_{3}=u_{4}=\textsc{False}$, then we would have the solution to (DRP) shown in Table 4.
$f_{1}$ $f_{2}$ $f_{3}$ $f_{4}$ $f_{5}$ $f_{6}$ $f_{7}$
$x_{1}$ 0 0 0 1 1 0 0
$x_{2}$ 0 0 1 1 1 0 0
$x_{3}$ 0 1 0 1 1 0 0
$x_{4}$ 1 0 1 1 1 0 0
$x_{5}$ 1 1 0 1 1 0 0
$x_{6}$ 1 1 1 1 1 0 0
$x_{7}$ 0 0 1 1 0 1 0
$x_{8}$ 0 1 1 0 0 1 0
$x_{9}$ 0 1 1 1 0 1 0
$x_{10}$ 1 0 1 1 0 1 0
$x_{11}$ 1 1 1 0 0 1 0
$x_{12}$ 1 1 1 1 0 1 0
$x_{13}$ 1 0 0 1 0 0 1
$x_{14}$ 1 0 1 0 0 0 1
$x_{15}$ 1 0 1 1 0 0 1
$x_{16}$ 1 1 0 0 0 0 1
$x_{17}$ 1 1 0 1 0 0 1
$x_{18}$ 1 1 1 0 0 0 1
$x_{19}$ 1 0 0 0 1 1 1
Table 4: Solution to (DRP) constructed from 3-SAT solution. The entries in bold are arbitrary. The italicized entries encode the solution of the corresponding 3-SAT problem.
Claim A.2.
If (DRP) is feasible then the 3-SAT instance is a YES instance.
Proof.
If (DRP) is feasible, then there exists one example which has features $V+1,\ldots,V+|L|$ equal to 1. This comes from the rightmost node of the auxiliary tree.
Without loss of generality, assume that such example is $x_{6|L|+1}$ (e.g., $x_{19}$ in Table 4).
Then in each of the trees $l\in L$, $x_{6|L|+1}$ must have gone to the right branch at the root. In this case, we know that $x_{6|L|+1}$ must fall into one of the leaves of the perfect binary tree that corresponds to a truth assignment that makes the clause $l$ satisfied.
A solution to 3-SAT can then be constructed by looking at the first $V$ components of $x_{6|L|+1}$. ∎
Appendix B Mixed-Integer Linear Programming Formulation
@@
B.1 Model Formulation (Without Bagging)
@@
B.2 Empirical Evaluation
The results are reported in Figure 4, and their run times are compared in Table 5. Note that the results for the CP model are those presented in Figure 2(a), repeated here to ease comparison. Comparing the different curves (which correspond to different maximum depth constraints) between Figures 4(a) and 4(b), we see that both approaches successfully solve the dataset reconstruction problem on COMPAS without the use of bagging to train the target random forests. Intuitively, the two feasibility models encode the same information, and define the same set of feasible reconstructions. Because they use different techniques to represent and explore it, they may end up with different reconstructions, but there is no a priori reason for one to outperform the other systematically, and as observed in our experiments, their reconstruction performances are generally similar.
@@ B.2 Empirical Evaluation
Figure 4: Average reconstruction error as a function of the number of trees $\lvert\mathcal{T}\rvert$ within the attacked forest $\mathcal{T}$, for different maximum depth values $d_{max}$ and for the random baseline. For the experiments on the COMPAS dataset, not using bagging, we report the results obtained using either the CP model (Section 6) or the MILP one (Section B.1).
@@ Appendix C Implementation Details for the CP Model
if $\textbf{if }\sum\limits_{c\in\mathcal{C}}y_{tvkc}\geq 1\textbf{ then}\quad\left(\bigwedge\limits_{i\in\Phi_{v}^{+}}{x_{ki}=1}\right)\land\left(\bigwedge\limits_{i\in\Phi_{v}^{-}}{x_{ki}=0}\right)$
@@
  • $\forall t\in\mathcal{T},\forall k\in\{1..N\},\forall v\in\mathcal{V}_{t}^{L}:$ if $\textbf{if }w_{tvk}=1\textbf{ then}\quad\left(\bigwedge\limits_{i\in\Phi_{v}^{+}}{x_{ki}=1}\right)\land\left(\bigwedge\limits_{i\in\Phi_{v}^{-}}{x_{ki}=0}\right)$